While buying a new car, Rachel made a down payment of $700 and agreed to make month-end payments of $320 for the next 5 years and 7 months. She was charged an interest rate of 2% compounded semi-annually for the entire term.

a. What was the purchase price of the car?

Round to the nearest cent

b. What was the total amount of interest paid over the term?
Answer #1

1.
Rate compounded monthly=(((1+2%/2)^(2/12)-1)*12)=1.9917172%

Purchase price of the car=700+320/(1.9917172%/12)*(1-1/(1+1.9917172%/12)^(12*5+7))
=20974.97513

2.
=320*(12*5+7)-(20974.97513-700)
=1165.02487